I really need your help on this problem. I do not understand it.
Find the area of the quadrilateral with vertices
A(−1, 2), B(4, 6), C(10, 6), and D(5, 2).

Answer by ewatrrr(24785)

 
Hi,

Find the area of the quadrilateral with vertices
A(−1, 2), B(4, 6), C(10, 6), and D(5, 2).
A = b*h = 6*4 = 24 units squared